### v1

@@ Abstract
High-dimensional tensor-valued data have recently gained attention from researchers in economics and finance. We consider the estimation and inference of high-dimensional tensor factor models, where each dimension of the tensor diverges. Our focus is on a factor model that admits CP-type tensor decomposition, which allows for non-orthogonal loading vectors. Based on the contemporary covariance matrix, we propose an iterative simultaneous projection estimation method. Our estimator is robust to weak dependence among factors and weak correlation across different dimensions in the idiosyncratic shocks. We establish an inferential theory, demonstrating both consistency and asymptotic normality under relaxed assumptions. Within a unified framework, we consider two eigenvalue ratio-based estimators for the number of factors in a tensor factor model and justify their consistency. Through a simulation study and two empirical applications featuring sorted portfolios and international trade flows, we illustrate the advantages of our proposed estimator over existing methodologies in the literature.
@@ 1 Introduction
Statistical methods and economic applications for the high-dimensional tensor factor analysis are still in their early stages of development. As in the classical panel setting, tensor factor models typically assume low-rank structures, with Canonical Polyadic (CP) and Tucker structures being the most common choices (see, e.g., Kolda and Bader, (2009)). Recent studies have explored various estimation approaches and extensions. For example, working with Tucker decomposition, Chen et al., (2022) considered two estimators based on the autocovariance matrices, while Han et al., 2022a extended these methods using an iterative procedure with the matrix unfolding mechanism. Chen and Fan, (2023) proposed an estimation method called $\alpha$-PCA that preserves the matrix structure and aggregates mean and contemporary covariance through a hyper-parameter $\alpha$. Chen and Lam, (2024) introduced a pre-averaging technique for the Tucker tensor factor model that significantly enhances the model’s inherent signal strength under certain conditions. Chen et al., (2024) introduced a semiparametric tensor factor model leveraging mode-wise covariates. In the context of CP decomposition, Han et al., (2023) proposed an iterative simultaneous orthogonalization algorithm with warm-start initialization, while Babii et al., (2023) employed tensor principal component analysis (TPCA), assuming orthogonal factor loadings. Chang et al., (2023) developed estimation procedure based on a generalized eigenanalysis constructed from the serial dependence structure of the underlying process.
In this paper, we focus on a tensor factor model with a CP low-rank structure due to its parsimonious features. We propose an iterative projection estimation based on contemporary covariance rather than autocovariance matrices. As highlighted by Chen and Fan, (2023), autocovariance-based methods rely on the assumption of non-zero autocovariances among individual factors, limiting their effectiveness in scenarios with serially independent factors or weak autocorrelations in tensor data. We develop inferential theory, establishing consistency, convergence rates, and limiting distributions under relaxed assumptions. Additionally, we extend the eigenvalue ratio-based estimator (Ahn and Horenstein, (2013)) for latent dimensions to tensor factor models and show estimation consistency.
@@
The remaining sections of this paper are organized as follows. Section 2 introduces the high-dimensional tensor factor model with a CP low rank structure allowing for nonorthogonal loading vectors. In Section 3, we present an iterative projection estimation procedure and two generalized eigenvalue ratio-based estimators for the number of the latent factors. Section 4 establishes the consistency and limiting distributions of the estimated loading vectors. We assess the finite sample performance through simulation in Section 5 and provide two empirical applications in Section 6. Finally, Section 7 concludes the paper with all mathematical proofs included in the Appendix.
@@ 1.1 Notations and preliminaries
Let $\|x\|_{q}=(x_{1}^{q}+...+x_{p}^{q})^{1/q}$, $q\geq 1$, for any vector $x=(x_{1},...,x_{p})^{\top}$. We employ the following matrix norms: matrix spectral norm $\|M\|_{2}=\underset{\|x\|_{2}=1,\|y\|_{2}=1}{\max}\|x^{\top}My\|_{2}=\sigma_{1}(M)$, where $\sigma_{1}(M)$ is the largest singular value of $M$. For two sequences of real numbers $\{a_{n}\}$ and $\{b_{n}\}$, we write $a_{n}\asymp b_{n}$ if there exists a constant $C$ such that $|a_{n}|\leq C|b_{n}|$ holds for all sufficiently large $n$, and $a_{n}\lesssim b_{n}$ if there exists a constant $C$ such that $a_{n}\leq Cb_{n}$.
@@
The mode-$k$ matricization of a tensor ${\cal A}\in\mathbb{R}^{d_{1}\times\cdots\times d_{K}}$ is denoted as $\hbox{\rm mat}_{k}({\cal A})\in\mathbb{R}^{d_{k}\times d_{-k}}$, where $d=\prod_{j=1}^{K}d_{j}$ and $d_{-k}=d/d_{k}=\prod_{j=1,j\neq k}^{K}d_{j}$. It is obtained by setting the $k$-th tensor mode as its rows and collapsing all the others into its columns. And the vectorization of the matrix/tensor ${\cal A}$ is denoted as ${\rm{vec}}({\cal A})\in\mathbb{R}^{d}$. Note that $\hbox{\rm mat}_{k}(\mbox{vec}({\cal A}))=\hbox{\rm mat}_{k}({\cal A})$.
@@ 2 Model
We consider a tensor-valued time series $\mathcal{Y}_{t}\in\mathbb{R}^{d_{1}\times d_{2}\times\cdots\times d_{K}}$, where $1\leq t\leq T$. Our focus is on a tensor factor model with a CP low-rank structure:
$\mathcal{Y}_{t}=\sum_{i=1}^{r}f_{it}\left(a_{i1}\otimes a_{i2}\cdots\otimes a_{iK}\right)+\mathcal{E}_{t},\quad t\leq T,$ (1)
where $\otimes$ denotes the tensor product, $f_{it}$ is a one-dimensional latent factor, $a_{ik}$ denotes the $d_{k}$-dimensional loading vector, which needs not to be orthogonal. Unlike Han et al., (2023), we permit arbitrary correlation structures among individual factors. Without loss of generality, we assume $\|a_{ik}\|_{2}=1$, for all $1\leq i\leq r$ and $1\leq k\leq K$. The noise tensor $\mathcal{E}_{t}$ is assumed to be uncorrelated with the latent factors but may exhibit weak correlations across different dimensions. The rank $r$ may either be fixed or divergent.
When $K=1$, $\mathcal{Y}_{t}$ reduces to a vector, and model (1) becomes the classical factor model, extensively studied in the literature (Bai and Ng, (2002) and Stock and Watson, (2002)). For $K>1$, an alternative approach is to vectorize data:
@@
where $\mbox{vec}({\cal Y}_{t})\in\mathbb{R}^{d}$ with $d=d_{1}d_{2}\cdots d_{K}$ and $F_{t}=\left(f_{1t},f_{2t},\cdots,f_{rt}\right)^{\top}\in\mathbb{R}^{r}$. However, this method ignores the tensor structure of the data and hence substantially increases the number of parameters in the loading matrices from $(d_{1}+d_{2}+\cdots+d_{K})r$ in the tensor case to $(d_{1}d_{2}\cdots d_{K})r$ in the stacked vector version. Our proposed approach, modeling $\mbox{vec}({\cal Y}_{t})$ as $AF_{t}+\mbox{vec}({\cal E}_{t})$, $A=(a_{1},\cdots,a_{r})$ and $a_{i}=\mbox{vec}(a_{i1}\otimes a_{i2}\otimes\cdots\otimes a_{iK})$, within our specific framework, yields improved convergence rates due to its unique structure.
Consider an illustrative example of international trade flows, detailed in Section 6. The observed $\mathcal{Y}_{t}$ forms a square matrix, where $d_{1}=d_{2}=n$ and $K=2$. Each entry $\mathcal{Y}_{t,ij}$ of $\mathcal{Y}_{t}$, with $i,j=1,2,\cdots,n$, represents the volume of trade flow from country $i$ to country $j$ at time $t$. Thus, the $i$th row represents data where country $i$ is the exporter, while the $j$th column represents data where country $j$ is the importer. Figure 1 shows a time series plot of $\mathcal{Y}_{t}$ for G7 countries excluding EU spanning from January 2008 to December 2014. Model (1) identifies $r$ latent factors, analogous to $r$ trading hubs. Each country exports to these hubs with certain distributions (determined by the loading matrix $A_{1}$) and imports from them likewise (determined by the loading matrix $A_{2}$). The element $a_{1,il}$ of $A_{1}$, where $i=1,\cdots,r$ and $l=1,\cdots,n$, represents the export contribution of country $l$ to trading hub $i$. Similarly, the entry $a_{2,jm}$ of $A_{2}$, where $j=1,\cdots,r$ and $m=1,\cdots,n$, can be interpreted as the import contribution of country $m$ to trading hub $j$. We allow the number of trading hubs $r$ to increase with the increase of $n$ and $T$.
Figure 1: Time series plots of the value of goods traded among G7 countriesU
  • Notes: (1) sample period: January 2008- December 2014. (2) The plots only show the patterns of the time series while the magnitudes are not comparable between plots because the ranges of the y-axis are different.
@@
where the core tensor ${\cal F}_{t}\in\mathbb{R}^{r_{1}\times\cdots\times r_{K}}$ is the latent factor process in a tensor form, and $A_{i}$’s are $d_{i}\times r_{i}$ loading matrices. As discussed in Babii et al., (2023) and Han et al., (2023), unlike the CP decomposition, the Tucker decomposition is generally non-unique, leading to significant identification issues. Consequently, estimation results from model (3) may exhibit ambiguity, undermining meaningful discussions of individual factors (Stock and Watson, (2002)). In contrast, the CP tensor factor model (1) yields a unique set of one-dimensional latent factors, which serve as natural inputs for diffusion index forecasts and factor-augmented regressions (Bai and Ng, (2006)). We regard the CP tensor factor as a more parsimonious yet flexible and effective alternative. Further comparison of the performance of these two tensor factor models will be presented in Section 6.

### v2

@@ Abstract
High-dimensional tensor-valued data have recently gained attention from researchers in economics and finance. We consider the estimation and inference of high-dimensional tensor factor models, where each dimension of the tensor diverges. Our focus is on a factor model that admits CP-type tensor decomposition, which allows for non-orthogonal loading vectors. Based on the contemporary covariance matrix, we propose an iterative simultaneous projection estimation method. Our estimator is robust to weak dependence among factors and weak correlation across different dimensions in the idiosyncratic shocks. We establish an inferential theory, demonstrating both consistency and asymptotic normality under relaxed assumptions. Within a unified framework, we consider two eigenvalue ratio-based estimators for the number of factors in a tensor factor model and justify their consistency. Simulation studies confirm the theoretical results and an empirical application to sorted portfolios reveals three important factors: a market factor, a long-short factor, and a volatility factor.
@@ 1 Introduction
Statistical methods and economic applications for high-dimensional tensor factor analysis are still in their early stages of development. As in the classical panel setting, tensor factor models typically assume low-rank structures, with Canonical Polyadic (CP) and Tucker structures being the most common choices (see, e.g., Kolda and Bader, (2009)). Recent studies have explored various estimation approaches and extensions. For example, working with Tucker decomposition, Chen et al., (2022) consider two estimators based on the autocovariance matrices, while Han et al., 2022a extend these methods using an iterative procedure with the matrix unfolding mechanism. Chen and Fan, (2023) propose an $\alpha$-PCA method that preserves the matrix structure and aggregates mean and contemporary covariance through a hyper-parameter $\alpha$. Chen and Lam, (2024) introduce a pre-averaging technique for the Tucker tensor factor model that significantly enhances the model’s inherent signal strength under certain conditions. In parallel, Lettau, (2024) provides a pioneering empirical application of Tucker decomposition to a three-dimensional panel of double-sorted portfolios, showing substantial improvements over Fama–French factors and traditional PCA. In the context of CP decomposition, Han et al., (2024) propose an iterative simultaneous orthogonalization algorithm with warm-start initialization, while Babii et al., (2022) employ tensor principal component analysis (TPCA), assuming orthogonal factor loadings. Chang et al., (2023) develop an estimation procedure based on a generalized eigenanalysis (GE) constructed from the serial dependence structure of the underlying process.
In this paper, we focus on a tensor factor model with a CP low-rank structure due to its economic relevance and parsimonious features. We propose an iterative projection estimation based on the contemporary covariance and develop new inferential theories. Our contributions advance the existing literature in several ways:
First, unlike Han et al., (2024), which rely on lagged autocovariances, we use the contemporary covariance, which captures the full variance structure of the data. This is crucial in applications where serial dependence is weak, such as asset returns, where the efficient market hypothesis implies little autocorrelation. Our method significantly broadens the applicability of CP tensor factor models.
Second, inspired by techniques in Anandkumar et al., 2014b and Sun et al., (2017) with noisy CP decomposition, we introduce an initialization strategy via randomized composite PCA (RC-PCA), which accommodates closely spaced eigenvalues. While those works consider CP decomposition of a single tensor, we establish performance bounds in the presence of noise and latent factor randomness—substantially complicating the theoretical analysis.
Third, we provide the first asymptotic normality results for estimated loading vectors in CP tensor factor models. This result enables valid statistical inference, such as confidence interval construction and hypothesis testing, tools that are essential for empirical applications [^2] The importance of developing inferential theory in factor models is well established in the econometric literature; for instance, Bai, (2003) laid the foundation for inference in high-dimensional factor models and has had a lasting impact. Our work brings similar inferential ideas to the tensor setting, filling an important gap in the existing literature.. Our inferential framework is based on the asymptotic distribution of singular vectors and differs fundamentally from the technical tools used in both vector factor models (Bai, (2003)) and Tucker tensor models (Chen and Fan, (2023), Yu et al., (2022)). Our approach can be extended to inference in many other matrix and tensor problems, such as the inference of factor loading vectors using lagged autocovariance matrices, as considered in Lam and Yao, (2012); Chang et al., (2023).
Fourth, we extend the eigenvalue ratio criterion—previously used in Tucker models (e.g., Han et al., 2022b )—to the CP tensor factor setting and establish its consistency. To our knowledge, this is the first work to do so.
Finally, our empirical analysis using characteristic-sorted portfolios demonstrates the practical value of our method. The proposed CP factor model outperforms benchmark approaches in out-of-sample forecasts. The factors extracted from the CP low-rank structure yield smaller cross-sectional pricing errors than classical factor models in the literature. Moreover, the estimated loadings and factors are interpretable without post-hoc rotation, revealing economically meaningful factors: a market factor, a long-short factor, and a volatility factor.
@@
The remaining sections of this paper are organized as follows. Section 2 introduces the high-dimensional tensor factor model with a CP low rank structure allowing for non-orthogonal loading vectors. In Section 3, we present an iterative projection estimation procedure and two generalized eigenvalue ratio-based estimators for the number of latent factors. Section 4 establishes inferential theory. We assess the finite sample performance through simulation in Section 5 and apply our method to characteristic portfolios in Section 6. Finally, Section 7 concludes the paper with all mathematical proofs and additional simulations included in the Appendix.
@@ 1.1 Notations and preliminaries
Let $[n]$ denote the set $\{1,2,\ldots,n\}$. Let $\|x\|_{q}=(x_{1}^{q}+...+x_{p}^{q})^{1/q}$, $q\geq 1$, for any vector $x=(x_{1},...,x_{p})^{\top}$. We employ the following matrix norms: matrix spectral norm $\|M\|_{2}=\underset{\|x\|_{2}=1,\|y\|_{2}=1}{\max}\|x^{\top}My\|_{2}=\sigma_{1}(M)$, where $\sigma_{1}(M)$ is the largest singular value of $M$. For two sequences of real numbers $\{a_{n}\}$ and $\{b_{n}\}$, we write $a_{n}\asymp b_{n}$ if there exists a constant $C$ such that $|a_{n}|\leq C|b_{n}|$ and $|a_{n}|\geq C|b_{n}|$ hold for all sufficiently large $n$, and $a_{n}\lesssim b_{n}$ if there exists a constant $C$ such that $a_{n}\leq Cb_{n}$.
@@
Similarly, for a matrix $V\in\mathbb{R}^{d_{k}\times d_{\ell}}$, define ${\cal A}\times_{\ell}\times_{k}V\in\mathbb{R}^{d_{1}\times\cdots\times d_{\ell-1}\times d_{\ell+1}\times\cdots\times d_{k-1}\times d_{k+1}\times\cdots\times d_{K}}$ as
$({\cal A}\times_{\ell}\times_{k}V)_{i_{1},...,i_{\ell-1},i_{\ell+1},...,i_{k-1},i_{k+1},...,i_{K}}=\sum_{i_{\ell}=1}^{d_{\ell}}\sum_{i_{k}=1}^{d_{k}}{\cal A}_{i_{1},i_{2},...,i_{K}}V_{i_{\ell},i_{k}}.$
The mode-$k$ matricization of a tensor ${\cal A}\in\mathbb{R}^{d_{1}\times\cdots\times d_{K}}$ is denoted as $\hbox{\rm mat}_{k}({\cal A})\in\mathbb{R}^{d_{k}\times d_{-k}}$, where $d=\prod_{j=1}^{K}d_{j}$ and $d_{-k}=d/d_{k}=\prod_{j=1,j\neq k}^{K}d_{j}$. It is obtained by setting the $k$-th tensor mode as its rows and collapsing all the others into its columns. And the vectorization of the matrix/tensor $\cal A$} is denoted as ${\rm{vec}}({\cal A})\in\mathbb{R}^{d}$. With a slight abuse of notation, we still define $\hbox{\rm mat}_{k}(\mbox{vec}({\cal A}))=\hbox{\rm mat}_{k}({\cal A})$. For nonempty $J\subseteq[K]$, $\text{mat}_{J}({\cal A})$ is the mode $J$ matrix unfolding which maps $\cal A$} to $d_{J}\times d_{-J}$ matrix with $d_{J}=\prod_{j\in J}d_{j}$ and $d_{-J}=d/d_{J}$, e.g. $\text{mat}_{\{1,2\}}({\cal A})=\text{mat}_{3}^{\top}({\cal A})$ for $K=3$.
@@ 2 Model
We consider a tensor-valued time series $\mathcal{Y}_{t}\in\mathbb{R}^{d_{1}\times d_{2}\times\cdots\times d_{K}}$, where $K\geq 2$ and $1\leq t\leq T$ [^3] When $K=1$, $\mathcal{Y}_{t}$ reduces to a vector, and model (1) becomes the classical factor model extensively studied in the literature (Bai and Ng, (2002) and Stock and Watson, (2002)). The identification condition discussed in Remark 2.1 does not apply in this case, and hence we assume $K\geq 2$. . Our focus is on a tensor factor model with a CP low-rank structure:
$\mathcal{Y}_{t}=\sum_{i=1}^{r}f_{it}\left(\Gamma_{i1}\otimes\Gamma_{i2}\cdots\otimes\Gamma_{iK}\right)+\mathcal{E}_{t}=\sum_{i=1}^{r}w_{i}f_{it}\left(a_{i1}\otimes a_{i2}\cdots\otimes a_{iK}\right)+\mathcal{E}_{t},\quad t\leq T,$ (1)
where $\otimes$ denotes the tensor product, $f_{it}$ is a one-dimensional latent factor, $\Gamma_{ik}$ denotes the $d_{k}$-dimensional loading vector, which needs not to be orthogonal [^4] Our framework naturally accommodates orthogonal loadings or factors as a special case, although our identification condition does not require orthogonal loadings or factors (Kolda and Bader,, 2009).. Without loss of generality and to ensure identifiability, we assume that $\mathbb{E}f_{it}^{2}=1$ and normalize the factor loadings $\Gamma_{ik}$ so that $a_{ik}=\Gamma_{ik}/\|\Gamma_{ik}\|_{2}$, for all $1\leq i\leq r$ and $1\leq k\leq K$. Consequently, all factor strengths are captured by $w_{i}$ with $w_{i}=\prod_{k=1}^{K}\|\Gamma_{ik}\|_{2}$. In the strong factor model case, $\|\Gamma_{ik}\|_{2}\asymp d_{k}^{1/2}$, which implies that $w_{i}\asymp\sqrt{d_{1}d_{2}\cdots d_{K}}$. Unlike the uncorrelated factors assumed in Han et al., (2024), we allow for moderately strong correlation structures among the individual factors. The noise tensor $\mathcal{E}_{t}$ is assumed to be uncorrelated with the latent factors but may exhibit weak correlations across different dimensions. The rank $r$ may either be fixed or divergent.
A natural alternative approach to extracting common factors is to vectorize the data:
@@
where $\mbox{vec}({\cal Y}_{t})\in\mathbb{R}^{d}$ with $d=d_{1}d_{2}\cdots d_{K}$ and $F_{t}=\left(f_{1t},f_{2t},\cdots,f_{rt}\right)^{\top}\in\mathbb{R}^{r}$. However, this method ignores the tensor structure of the data and hence substantially increases the number of parameters in the loading matrices from $(d_{1}+d_{2}+\cdots+d_{K})r$ in the tensor case to $(d_{1}d_{2}\cdots d_{K})r$ in the stacked vector version. In contrast, our proposed model preserves the tensor structure by modeling $\mbox{vec}({\cal Y}_{t})$ as $AWF_{t}+\mbox{vec}({\cal E}_{t})$, $W=\text{diag}(w_{1},...,w_{r}),A=(a_{1},\cdots,a_{r})$ and $a_{i}=\mbox{vec}(a_{i1}\otimes a_{i2}\otimes\cdots\otimes a_{iK})$. This structure not only reduces the parameter dimension but also leads to improved convergence rates due to its parsimonious structure.
Consider an illustrative example of sorted portfolios, detailed in Section 6. The observed $\mathcal{Y}_{t}$ is represented as a matrix, where $d_{1}$ is the number of characteristic-sorted portfolios, $d_{2}=10$, and $K=2$. Each entry $\mathcal{Y}_{t,jl}$ corresponds to the excess return of the $l^{\text{th}}$-decile of the $j^{\text{th}}$ characteristic at time $t=1,\ldots,T$. Figure 1 shows a time series plot of $\mathcal{Y}_{t}$ for ten characteristics from January 1990 to December 2022. Model (1) identifies $r$ latent factors, which can be interpreted as systematic risk factors. The element $a_{i1,j}$ of the loading vector $a_{i1}$, where $i=1,\cdots,r$ and $j=1,\cdots,d_{1}$, captures the heterogeneous exposure of the $j^{\text{th}}$ characteristic to the $i^{\text{th}}$ risk factor. Similarly, the entry $a_{i2,l}$ of the loading vector $a_{i2}$, where $i=1,\cdots,r$ and $l=1,\cdots,10$, determines the exposure of the $l^{\text{th}}$ decile to the $i^{\text{th}}$ risk factor. We allow the number of risk factors $r$ to increase with the dimensions $d_{1}$, $d_{2}$ and the sample size $T$.
Figure 1: Time series plots of characteristic decile portfolio returns
  • Note: sample period from January 1990 to December 2022.
@@
where the core tensor ${\cal F}_{t}\in\mathbb{R}^{r_{1}\times\cdots\times r_{K}}$ is the latent factor process in a tensor form, and $A_{i}$’s are $d_{i}\times r_{i}$ loading matrices. As discussed in Babii et al., (2022) and Han et al., (2024), unlike the CP decomposition, the Tucker decomposition is generally non-unique, leading to identification difficulties: even with the usual identification restrictions, only the column spaces of loading matrices are identified. Consequently, estimation results from model (3) may exhibit ambiguity, undermining meaningful discussions of individual factors [^5] For classical factor models, Stock and Watson, (2002) point out, “because the factors are identified only up to a $k\times k$ matrix, detailed discussion of the individual factors is unwarranted”. The same comment applies to Tucker tensor factor models.. In contrast, the CP tensor factor model (1) yields $r$ scalar factors, aligning with the conventional wisdom in economic applications regarding common factors. This set of one-dimensional latent factors serves as natural inputs for diffusion index forecasting and factor-augmented regressions [^6] A prominent example is the modeling of global yield curve dynamics. Diebold et al., (2008) show that the generalized Nelson-Siegel model accurately captures the dynamics of yield curves and delivers strong predictive performance. Their model assumes that global yields follow a linear factor structure with interpretable loadings. This specification fits naturally within our CP factor model framework, where $f_{t}$ captures global factors, and $a_{i1}$ and $a_{i2}$ represent country-specific and maturity-specific loadings, respectively.. We regard the CP tensor factor as a more parsimonious yet flexible and economically relevant alternative. Further comparison of the performance of these two tensor factor models will be presented in Section 6.
Remark 2.1 (Identifiability Condition).
By incorporating time, we may stack ${\cal Y}_{t}$ into an order-$(K+1)$ tensor ${\cal Y}\in\mathbb{R}^{d_{1}\times\cdots\times d_{K}\times T}$, with time $t$ as the $(K+1)$-th mode, referred to as the time-mode. Subsequently, model (1) can be reformulated as
${\cal Y}=\sum_{i=1}^{r}w_{i}a_{i1}\otimes a_{i2}\otimes\cdots\otimes a_{iK}\otimes\boldsymbol{f}_{i}+{\cal E},$
[…]
The identifiability condition is sufficient but not necessary—it is conservative. Identification can still hold even when it fails, but the guarantee is lost. It will definitely fail (and identification will typically be weak or lost) when a mode has very low rank, e.g., if some loading vectors are constant or proportional so that many columns are nearly collinear. For instance, if one mode effectively has rank =1=1 (e.g., several components share the same constant loading vector), then the sum of ranks can’t clear the threshold unless the rank is trivial; in practice, this allows components to be mixed within the low-rank subspace, recreating a rotation-like ambiguity. More broadly, near-violations (high collinearity, duplicate columns, extremely unbalanced ranks/dimensions) lead to weak identification and numerical instability even if the identifiability condition barely holds.